OpenStudy (anonymous):

how do I make jonathan te subject of an equation if molly and louise are twins(his sisters) and he is four years younger? their total ages are 17 years

OpenStudy (anonymous):

let molly n louise age(twin)=x (both have same age) jonathan age=x-4 ( Mr. J is 4 year younger) Now, x+x+x-4=17 (Total age) That's it.

jhonyy9 (jhonyy9):

let x+x=2x represent the age of Molly and Louise so and y =x-4 is the age of Jonathan and you know again that their total ages are 17 years so this mean that 2x+y=17 and you know again that y =x-4 so 2x+x-4=17 so 3x=21 x=7 y=x-4 y=3 hope this is understandably and this will help you

OpenStudy (aum):

"make Jonathan the subject of the equation" Let x represent Jonathan's age. He is 4 years younger than his sisters. That means the sisters are 4 years older than him. Therefore each twin sister must be (x + 4) years old. Sum of their ages = x + (x+4) + (x+4) = 17 or 3x + 8 = 17 is the equation. If you are asked to solve the equation: 3x + 8 = 17 subtract 8 from both sides 3x = 17 - 8 3x = 9 divide both sides by 3 x = 3. So Jonathan is 3 years old and his sisters are each (x+4 = 3+4 =) 7 years old.